OpenStudy (lilkg77):

HELP!!!!!!! A certain type of parasite measures .0012 mm on its first day of life. It then triples in length each day after that. How many days will it take for the parasite to grow to be 150 mm? A. About 5.1 days B. About 10.7 days C. About 49.9 days D. About 41,666.7 days

OpenStudy (lilkg77):

HELP PLZ!

OpenStudy (priyar):

let the no. of. days be "x"..

OpenStudy (priyar):

can u frame an eq for this problem?

OpenStudy (lilkg77):

If you frame one for me i might be able to solve it

OpenStudy (priyar):

ok then..on the first day it is 0.0012..on the second day it is 3*0.0012..on the third day it is 3*(3*0.0012)..and so on

OpenStudy (priyar):

do u notice a pattern here with the no of days..?

OpenStudy (mathmale):

So there is an increase of ( what ) from each term to the next? Phrased differently, is there a factor that will take you from the parasite's length one day to its length on the next day?

OpenStudy (priyar):

when x=1 0.0012 x=2 3^1(0.0012) x=3 3^2(0.0012)

OpenStudy (mathmale):

Write a model for the length of the parasite. L=(?) (?) ^ (?)

OpenStudy (lilkg77):

That answer is definitely D right? @priyar @mathmale

OpenStudy (priyar):

how can u tell?

OpenStudy (priyar):

if the day starts from 0 instead of one the eq will be of the form...150= 3^x..now find x

OpenStudy (mathmale):

By testing your own equation. Does it correctly predict length after a given number of days?

OpenStudy (priyar):

sorry typing error..it is 150=0.0012(3^x)

OpenStudy (mathmale):

Assuming that this model is correct, how would you solve for time, x? Hint: What is the inverse function of the exponential function 3^x?

OpenStudy (priyar):

D is wrong @lilkg77 ..try solving for x...first divide by 0.0012 on both sides..what do u get?

OpenStudy (priyar):

sorry the eq is: 150=0.0012*(x^3) coz in the q it says triples and not three times..

OpenStudy (priyar):

now divide by 0.0012 on both sides..

OpenStudy (priyar):

what do u get @lilkg77 ? pls respond..

OpenStudy (priyar):

??

OpenStudy (priyar):

anyway u get 125000=x^3..so what is x?

OpenStudy (priyar):

ok..x=50 since 50*50*50=125000..so what will be the answer @lilkg77 ??

OpenStudy (lilkg77):

Sorry @priyar i was in using the bathrooom

OpenStudy (lilkg77):

The answer would be C @priyar

OpenStudy (misty1212):

HI!!

OpenStudy (priyar):

its ok..and u are correct! did u understand?

OpenStudy (lilkg77):

Yeah thanks! @priyar

OpenStudy (misty1212):

i do not think it is C

OpenStudy (lilkg77):

? @misty1212

OpenStudy (misty1212):

it is not \[150=.0012\times x^3\] it is \[150=.0012\times 3^x\]

OpenStudy (misty1212):

you need logs to solve this

OpenStudy (misty1212):

divide by .0012 is the first step

OpenStudy (priyar):

@misty1212 even i thought so at first!(scroll up) but they have given "triples" and not "three times"

OpenStudy (misty1212):

triples yes

OpenStudy (misty1212):

1, 3, 9, 27, 81,...

OpenStudy (misty1212):

\[125,000=3^x\\ x=\frac{\ln(125000)}{\ln(3)}\] and a calculator

OpenStudy (misty1212):

or you can solve it directly using wolfram http://www.wolframalpha.com/input/?i=150%3D.0012*3%5Ex